A metal alloy weighing 10 lb. and containing 10% nickel is melted and mixed with 8 lb. of pure nickel. What percent of the resulting alloy is nickel?
Answer by lwsshak3(11628)

A metal alloy weighing 10 lb. and containing 10% nickel is melted and mixed with 8 lb. of pure nickel. What percent of the resulting alloy is nickel?
**
let x=percent of nickel the resulting alloy
..
10%(10)+100%(8)=x(18)
.1+8=18x
x=8.1/18=.45=45%